OpenStudy (calculusxy):

How many two digit numbers are divisible by either 3 or 5? @Mertsj

OpenStudy (anonymous):

how many two digit numbers are there?

OpenStudy (calculusxy):

99

OpenStudy (anonymous):

if the number of two digit numbers is x, then 9>x>100

OpenStudy (anonymous):

it is hence not 99

OpenStudy (anonymous):

49

OpenStudy (anonymous):

there are actually 90 two digit numbers, take 99-9

OpenStudy (anonymous):



OpenStudy (anonymous):

how many numbers from 10 to 99 are divisible by 5 or 3?

OpenStudy (anonymous):

It's not 3 AND 5.

OpenStudy (calculusxy):

The options are 42, 46, 47, 48, 53

OpenStudy (calculusxy):

So how do I figure this out?

OpenStudy (anonymous):

ok, so find the total number of number divisible by 3 OR 5

OpenStudy (anonymous):

what is it?

OpenStudy (mertsj):

The ones divisible by 5 are 10, 15, 20, 25...95 so there are 18 of those. The ones divisible by 3 are the ones in which the sum of the digits is a multiple of 3

OpenStudy (calculusxy):

90/3=30 90/5=18 When I add that I get 48. That answer is not correct according got the answer key.

OpenStudy (mertsj):

So that would be 12, 15, 18, 21, 24, 27...

OpenStudy (calculusxy):

@satellite73

myininaya (myininaya):

I would start by thinking multiplies of 3 and 5. \[3n_1 \text{ is all the multiplies of 3} \] \[5n_2 \text{ is all the multiplies of 5} \] We want \[12 \le 3n_1 \le 99 \text{ and } 20 \le 5n_2 \le 95 \] where \[n_1 \text{and} n_2 \text{ are integers }\] Solve both inequalities

OpenStudy (anonymous):

99/3+(100/5-1), did u get that?

OpenStudy (anonymous):

0 duhhh

OpenStudy (calculusxy):

@utopia1234 Why do you have to subtract 1 from it?

myininaya (myininaya):

I named the first multiple of 3 which is 12 that is a 2 digit noticed i started the second set at 20 instead of 15 because 15 was already in the first set

OpenStudy (anonymous):

33 + 19=52 (number of numbers divisible by 3 or 5 from 1 to 99)

myininaya (myininaya):

I got something less that than @utopia1234

OpenStudy (anonymous):

99/15 round down, u will get 6

OpenStudy (anonymous):

52-6 will be the number that that is divisible by 3 or 5 in the first 99 numbers

myininaya (myininaya):

Yeah! that is what I got.

OpenStudy (anonymous):

numbers divisible by 3 and 5 from 1 to 10 includes 3,6,9,5

OpenStudy (anonymous):

52-6-4 will give u 42 which is the answer

OpenStudy (mertsj):

10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95, 12, 18, 21, 24, 27, 33, 36, 39, 42, 48, 51, 54, 57, 63, 66, 69, 72, 78, 81, 84, 87, 93, 96, 99

OpenStudy (mertsj):

Count them.

OpenStudy (anonymous):

LOL ok heres the method, or the way i was thinking.

myininaya (myininaya):

Well how about this 12<=3n_1<=99 where n_1 is an integer solve for n_1 4<=n_1<=33 There are 33-4+1 numbers in this set Then there is also this set 20<5n_2<95 where n_2 is an integer solve for n_2 4<=n_2<=19 There are 19-4+1 numbers in this set. So the answer should be (33-4+1)+(19-4+1)

myininaya (myininaya):

I'm meaning the inequality above to be a set of integers not rationals and irrationals just integers

OpenStudy (anonymous):